# How to Train a Backdoor-Robust Model on a Poisoned Dataset without Auxiliary Data?

### v1

@@ Abstract
Backdoor attacks have attracted wide attention from academia and industry due to their great security threat to deep neural networks (DNN). Most of the existing methods propose to conduct backdoor attacks by poisoning the training dataset with different strategies, so it’s critical to identify the poisoned samples and then train a clean model on the unreliable dataset in the context of defending backdoor attacks. Although numerous backdoor countermeasure researches are proposed, their inherent weaknesses render them limited in practical scenarios, such as the requirement of enough clean samples, unstable defense performance under various attack conditions, poor defense performance against adaptive attacks, and so on.
Therefore, in this paper, we are committed to overcome the above limitations and propose a more practical backdoor defense method. Concretely, we first explore the inherent relationship between the potential perturbations and the backdoor trigger, and the theoretical analysis and experimental results demonstrate that the poisoned samples perform more robustness to perturbation than the clean ones. Then, based on our key explorations, we introduce AdvrBD, an Adversarial perturbation-based and robust Backdoor Defense framework, which can effectively identify the poisoned samples and train a clean model on the poisoned dataset. Constructively, our AdvrBD eliminates the requirement for any clean samples or knowledge about the poisoned dataset (e.g., poisoning ratio), which significantly improves the practicality in real-world scenarios. Finally, we conduct extensive experiments and evaluate the effectiveness and superiority of the proposed AdvrBD in defending against $8$ state-of-the-art attacks (including $3$ adaptive attacks) on $4$ datasets (Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100). The experimental results demonstrate that our AdvrBD can achieve the average ASR with nearly $0.00\%$ in defending against $8$ state-of-the-art backdoor attacks while maintaining high accuracy of the main task. Compared with the $7$ state-of-the-art backdoor defense methods, our AdvrBD also performs more prominently. Furthermore, the excellent performance on $3$ different model architectures and poisoning ratios also highlights the remarkable generalization capability of AdvrBD.
@@ I Introduction
Training deep learning models relies on a large amount of training data [1, 2, 3]. However, the training data are usually collected from the web or purchased from an untrustworthy third party, which leads to a risk of being poisoned or unreliable [4, 5]. For example, a company that wants to train a facial recognition model is likely to obtain poisoned data from a third party who aims to hijack the facial recognition model for profit [6, 7]. Therefore, the unreliable training dataset may bring many serious security threats to the model training side. One of the typical threats is the backdoor attack, which is usually conducted by poisoning the training data [8, 9]. In the backdoor attack, the attacker manipulates a few training samples by adding a specific trigger (e.g., a pixel patch or blending) and modifies the labels of these samples as a target class. The model trained on such a manipulated dataset can learn the main task while maintaining a backdoor function that can support the attacker in hijacking the model. The backdoored model maintains high accuracy on the normal samples but predicts the target class whenever the trigger pattern is attached to a test sample, as depicted in Fig.1. Backdoor attacks bring a significant security risk for model training owners as they are easily conducted and allow adversaries to control the trained model stealthily [10, 11]. For example, a backdoored traffic sign recognition model may misclassify a stop sign with the trigger patch as a no-tooting sign, which may cause a serious traffic accident. A face recognition model may misclassify an illegal user with the trigger as a valid user, which may pose a tremendous threat to the property of the valid user.
Figure 1: Overview of backdoor attacks.
@@
In recent years, backdoor attacks have attracted much attention from researchers due to their serious threat to model security. Numerous studies focus on how to identify the poisoned samples and train a clean model on the poisoned dataset [12, 13, 14, 15, 16]. However, the existing defense methods have some non-negligible limitations in practical applications. (1) Requirement for clean samples. Their defense effectiveness heavily relies on obtaining a substantial number of clean samples [14, 15, 16], which is impractical for defenders in certain specific real-world scenarios [17]. (2) Unstable defense performance. Their actual defense performance varies significantly depending on the choice of backdoor attack methods and poisoning ratios [18, 13], which is usually unknown to defenders in advance in practical applications. (3) Inadequate defense against adaptive attacks. They may exhibit poor resilience against adaptive attacks [13], e.g., transforming some poisoned samples.
@@
Therefore, in this paper, we aim to overcome the shortcomings mentioned above and propose a more practical and robust backdoor defense approach. To achieve these goals, there are three challenges that we need to address:
@@
To address the above challenges, we propose an Adversarial perturbation-based and robustness Backdoor Defense (AdvrBD) framework. Concretely, we first explore the intrinsic relationship between perturbation and backdoor through rich theoretical analysis and experimental verification. Based on the above exploration results, we roughly isolate the poisoned samples and clean samples by adding adversarial perturbation to the training data and then observing the change in the output probability distribution. Note that we just select some clean samples rather than precisely distinguish the poisoned samples in this step because of the certain overlap between the poisoned and the clean ones. Then, we train an enhanced backdoored model by unlearning the selected clean samples and relearning the remaining poisoned dataset. The poisoned samples and the clean samples can be precisely identified based on the enhanced backdoored model. Finally, we can train a clean model on the identified clean samples. To improve the performance of the clean model, we further fine-tune the clean model by relabeling the poisoned samples and adding them to the clean dataset. The main contributions are threefold:
  • Firstly, we discover an important observation that the poisoned samples present more robust to the potential perturbations than the clean samples. The theoretical analysis and experimental results also demonstrate our findings.
  • Then, we propose a novel and robust backdoor defense (namely AdvrBD) framework without any auxiliary clean dataset, which can resist various popular backdoor attacks and multiple adaptive attacks.
@@
  • Finally, extensive experimental results demonstrate that when faced with $8$ state-of-the-art backdoor attacks on $4$ datasets, our AdvrBD outperforms the $7$ state-of-the-art backdoor defense methods in terms of decreasing the attack success rate (average ASR nearly $0.00\%$) while only incurring negligible accuracy loss. Moreover, our AdvrBD has a satisfactory generalization on multiple model architectures and different poisoning ratios.
The remainder of the paper is as follows. In Section II, we introduce the related works. Then, we will show the preliminaries and the threat model in Section III and Section IV, respectively. In Section V, we present the design of the proposed backdoor defense strategy. The experimental results are shown and analyzed in Section VI. Finally, we summarize in Section VII.
II Related Works
@@
In this section, we review the relevant works of backdoor attacks and introduce the development of backdoor defense methods.
II-A Backdoor Attack Methods
@@
In recent years, many backdoor attack approaches have been proposed, including patch-based attacks, visibility of trigger attacks, label consistency, and so on. Gu et al.[9] proposed a typical backdoor attack method by adding a specific patch on the samples and modifying the corresponding label as the targeted label. This attack achieves high attack success and has a low-performance impact on the main task. Chen et al.[19] proposed a backdoor attack in which the trigger is blending background rather than a single pixel. This means that the trigger is hard for human beings to notice. Liu et al.[20] proposed a more stealthy backdoor attack that plants reflections as a trigger into the victim model. This attack can be resistant to many existing defense methods. Liu et al.[21] proposed a lightweight backdoor attack that just needs to inverse the neural network to generate a general trojan trigger and fine-tune some layers of the network to implant the trigger. However, all the above backdoor attacks require poisoning the label, which is easier to detect. Therefore, some researchers have also proposed clean-label backdoor attacks. For example, Shafahi et al.[22] proposed a targeted clean-label poisoning attack. This attack crafts poison images that collide with a target image in feature space, thus making it undistinguishable from a network. Because the attacker does not need to control the label, it is more stealthy to conduct a backdoor attack. Turner et al.[23] proposed a clean-label backdoor attack based on adversarial examples and GAN-generated data. The key feature of this attack is that the poisoned samples appear to be consistent with their label and thus seem clean even from human inspection. Chen et al.[7] proposed an Invisible Poisoning Attack (IPA), which is difficult to detect by existing defense methods. This attack not only employs highly stealthy poison training examples with the clean labels (perceptually similar to their clean samples), but also does not need to modify the labels. Li et al.[24] proposed two stealthy backdoor attacks in which the triggers are derived from the covert features. Compared with the existing backdoor attacks, the trigger patterns of this attack method are invisible to human eyes. Moreover, it is difficult to recover the backdoor trigger through the optimization algorithm. Zhu et al.[25] proposed a transferable clean-label poisoning attack in which the poison samples are fabricated to surround the targeted sample in feature space. Saha et al.[26] proposed a novel backdoor attack in which the poisoned samples are similar to the clean samples with the correct labels. [27] and [28] proposed defense-resistant backdoor attacks in outsourced cloud environment.
II-B Backdoor Defense Methods
@@
Although many backdoor defense methods have been proposed, they have some limitations (e.g., requiring an auxiliary clean dataset) used in practical scenarios. Different from the existing backdoor defense methods, we first explore the relationship between the perturbation and the backdoor. Then, based on the exploration results, we plan to propose a robust backdoor defense method that does not require an auxiliary clean dataset and has a stable performance against various backdoor attacks on different models and poisoning ratios.
III Preliminaries
In this section, we introduce the mainly related technologies in this paper.
III-A Backdoor Attack
A backdoor attack is to inject a trigger into a model by poisoning the training dataset. During the inference period, the backdoored model performs well on the original task but outputs specific attacker-chosen answers when the input contains a specific trigger. For more clarity, we formalize the most common backdoor attack method BadNets [9] as follows:
@@ III-A Backdoor Attack
where $\times$ is pixel-wise multiplication and $\boldsymbol{x}^{\prime}$ is the poisoned training sample. After modifying the training sample, the corresponding label is fixed with a target label. A successful backdoored model will maintain high accuracy on the main task while outputting the target label when the trigger $\boldsymbol{t}$ appears. After BadNets [9], many different attack variations have been proposed to enhance effectiveness and stealthiness. For example, Blend [19] and PhysicalBA [40] have designed more complex patterns. WaNet [41] has proposed a stealthier input-specific-trigger attack. SIG [42] has proposed a clean-label attack that is stealthier as it does not change the labels. To evade the existing defense method, some adaptive attacks (e.g., TaCT [37], AdaptiveBlend [43] and AdaptivePatch [43]) have also been proposed to improve the attack performance.
Figure 2: Threat Model of AdvrBD.
IV Threat Model
In this paper, we consider a practical scenario where a model owner aims to train a deep learning model based on the training dataset provided by an untrusted third party. However, the third party may poison the training dataset to hijack the model in the future. We assume that the model owner and the untrusted third party play the role of the defender and the attacker in the threat model, respectively. The threat model is shown in Fig.2. The exact goal, knowledge, and capabilities of the defender and the attacker are described in the following.
@@ IV Threat Model
Figure 3: Framework of AdvrBD.
V Design of AdvrBD
@@
V-A Overview of AdvrBD
In this paper, we focus on how to identify the poisoned samples and train a clean model on the poisoned dataset. It is not trivial to satisfy this goal since even a tiny number of poisoned samples can accomplish the backdoor insertion process. To tackle this problem, we explore the inherent relationship between the perturbation and the backdoor attack theoretically and experimentally at first. Based on the exploration results, we propose a three-stage backdoor defense method dubbed AdvrBD and the overall framework is shown in Fig.3. There are three main phases in the proposed AdvrBD. Firstly, we partition a tiny number of clean samples by using the difference between clean samples and poisoned samples in the face of perturbations. An enhanced backdoored model is trained by unlearning the chosen clean samples and relearning the remaining dataset. Then, we precisely distinguish the poisoned samples and clean samples with a high true positive rate based on the enhanced backdoored model. The samples that are misclassified by the enhanced backdoored model are defined as clean samples. The rest are the poisoned samples. A clean model can be trained based on the identified clean samples. Finally, we relabel the poisoned samples and combine them with the clean samples to obtain a completely clean dataset, which can be used to fine-tune the clean model to further improve the model performance. The design of AdvrBD is presented in detail as follows.
V-B Detailed Design of AdvrBD
@@
V-B1 Exploring the Relationship between Perturbation and Backdoor
To identify the poisoned samples accurately, we need to train an enhanced backdoored model by unlearning some clean samples. It’s vital to select a few clean samples without any auxiliary dataset and figure out how poisoned samples influence the prediction of backdoored models, through which we can design corresponding metrics to separate the clean samples and poisoned ones. We begin our exploration by reviewing the backdoor attack process first.
Let $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ be a clean training set and $C:\mathcal{X}\rightarrow\mathcal{Y}$ denotes the functionality of the target neural network. For each image $\boldsymbol{x}_{i}$ in $\mathcal{D}$, we have $\boldsymbol{x}_{i}\in\mathcal{X}=[0,1]^{C\times W\times H}$, and $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$ is the corresponding label, where $K$ is the number of label classes. To launch an attack, backdoor adversaries first need to poison the selected clean samples $\mathcal{D}_{p}$ with covert transformation $T(\cdot)$. Then the poisoned samples are mixed with clean ones before training a backdoored model, the process of which can be formalized as: $\mathcal{D}_{t}=\mathcal{D}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}={(x_{i}^{\prime},y_{t})|x^{\prime}=T(x),(x_{i},y_{i})\in\mathcal{D}_{p}}$. Many attacks have been proposed to make $T(\cdot)$ more stealthy and undetectable or make less assumption towards adversaries’ capability like poison ratio $pr=|\mathcal{D}_{p}|/|\mathcal{D}|$. Nevertheless, these adversaries hold the same goal to hijack the neural networks by training a malicious network:

### v2

# Mellivora Capensis: A Backdoor-Free Training Framework on the Poisoned Dataset without Auxiliary Data

@@ Abstract
The efficacy of deep learning models is profoundly influenced by the quality of their training data. Given the considerations of data diversity, data scale, and annotation expenses, model trainers frequently resort to sourcing and acquiring datasets from online repositories. Although economically pragmatic, this strategy exposes the models to substantial security vulnerabilities. Untrusted entities can clandestinely embed triggers within the dataset, facilitating the hijacking of the trained model on the poisoned dataset through backdoor attacks, which constitutes a grave security concern. Despite the proliferation of countermeasure research, their inherent limitations constrain their effectiveness in practical applications. These include the requirement for substantial quantities of clean samples, inconsistent defense performance across varying attack scenarios, and inadequate resilience against adaptive attacks, among others. Therefore, in this paper, we endeavor to address the challenges of backdoor attack countermeasures in real-world scenarios, thereby fortifying the security of training paradigm under the data-collection manner. Concretely, we first explore the inherent relationship between the potential perturbations and the backdoor trigger, and demonstrate the key observation that the poisoned samples perform more robustness to perturbation than the clean ones through the theoretical analysis and experiments. Then, based on our key explorations, we propose a robust and clean-data-free backdoor defense framework, namely Mellivora Capensis (MeCa), which enables the model trainer to train a clean model on the poisoned dataset. MeCa can effectively identify poisoned samples and train a clean model on the poisoned dataset and eliminate the requirement of clean samples or knowledge about the poisoned dataset (e.g., poisoning ratio). We conduct extensive experiments in defending against $8$ SOTA attacks (including $3$ adaptive attacks) on $4$ datasets (Imagnette, Tiny ImageNet, CIFAR10, and CIFAR100). The experimental results reveals that MeCa can achieve an average ASR with almost $0.00\%$ to defend against SOTA backdoor attacks while maintaining model availability, which outperforms $7$ SOTA backdoor defense methods. Furthermore, the excellent performance on $3$ different model architectures and poisoning ratios also highlights the remarkable generalization capability of MeCa.
@@ 1 Introduction
Training deep learning models relies on a large amount of training data [1, 2, 3, 4]. However, the training data are usually collected from the web or purchased from an untrustworthy third party, which leads to a risk of being poisoned or unreliable [5, 6]. For example, a company that wants to train a facial recognition model is likely to obtain poisoned data from a third party who aims to hijack the facial recognition model for profit [7, 8], as depicted in Fig.1. Therefore, the unreliable training dataset may bring many serious security threats to the model training side, which breaches the national regulations, such as Artificial Intelligence Act (AIA) [9], Digital Services Act (DSA) [10]. One of the typical threats is the backdoor attack, which is usually conducted by poisoning the training data [11, 12]. In the backdoor attack, the attacker manipulates a few training samples by adding a specific trigger (e.g., a pixel patch or blending) and modifies the labels of these samples as a target class. The backdoor model maintains high accuracy on the normal samples but predicts the target class whenever the trigger pattern is attached to a test sample, as depicted in Fig.1. Backdoor attacks bring a significant security risk for model training owners as they are easily conducted and allow adversaries to control the trained model stealthily [13, 14]. For example, a backdoored traffic sign recognition model may misclassify a stop sign with the trigger patch as a no-tooting sign, which may cause a serious traffic accident. Therefore, untrusted sources of datasets often serve as a breeding ground for backdoor attacks, posing a significant security threat to model trainers.
Regarding the third-party poisoned training dataset, both academia and industry have conducted extensive research, such as the detection and purification of backdoor samples in poisoned datasets [15, 16]. Although effective backdoor detection and purification are valuable, economically speaking, it is clearly more cost-effective and practical to train a clean model directly on a poisoned dataset if that were possible. Such paradigm simplifies data prep and reduces cleaning costs by enabling training on potentially tainted datasets, fostering robust models without extensive preprocessing. It’s crucial in resource-intensive data scenarios and could redefine data handling across fields, ensuring reliable, cost-effective outcomes [17], which implies we possess the capability to rebuild amidst the ruins.
Figure 1: Security threats on training the poisoned data.
@@
Consequently, another research direction in backdoor defense is to train clean models on poisoned datasets [18, 19, 20, 21, 22]. However, the existing methods have some non-negligible limitations in practical applications. (1) Requirement for clean samples. Their defense effectiveness is highly dependent on the obtaining of a substantial number of clean samples [20, 21, 22], which is impractical for defenders in certain specific real-world scenarios [23]. (2) Unstable defense performance. Their actual defense performance varies significantly depending on the choice of backdoor attack methods and poisoning ratios [24, 19], which is usually unknown to defenders in advance in practical applications. (3) Inadequate defense against adaptive attacks. They may exhibit poor resilience against adaptive attacks [19], e.g., transforming some poisoned samples.
Therefore, in this context, we ask the following research questions: {mdframed}[backgroundcolor=black!10,rightline=false,leftline=false,topline=false,bottomline=false,roundcorner=2mm] Is it feasible to train a clean model directly on a poisoned dataset, without any knowledge of the attacker’s capabilities (such as poisoning ratio) and without any auxiliary clean dataset? If so, how to ensure model availability?
@@
In this paper, we intend to overcome the shortcomings mentioned above and propose a more practical and robust backdoor defense approach. To achieve these goals, there are three challenges that we need to address:
@@
To address above challenges, we propose the Mellivora Capensis (named MeCa) framework, a robust and clean-data-free backdoor defense scheme based on adversarial perturbation, which enables the defender to train a clean model on a poisoned dataset. MeCa employs a three-step approach to achieve the objective. First, we delve into the fundamental relationship between perturbations and backdoors through comprehensive theoretical analysis and rigorous experimental validation. Then, we utilize adversarial perturbations to distinguish between clean samples and backdoor samples at a coarse granularity based on the above key observation about perturbations and backdoors. By unlearning identified clean samples and re-training on the remaining poisoned dataset, we refine a backdoor model that precisely distinguishes between poisoned and clean samples. Finally, we train a clean model on the identified pristine samples and enhance its performance through fine-tuning, incorporating relabeled poisoned samples into the clean dataset. The main contributions are threefold:
  • We explore the intrinsic differences between poisoned samples and clean samples, uncovering key observation about the differences in their perturbation manifestations. We conducted extensive theoretical analysis and experiments to substantiate our findings.
  • We propose a robust and clean-data-free backdoor defense framework, namely Mellivora Capensis (MeCa). MeCa allows defenders to train a clean model directly on poisoned data that maintains high accuracy on the primary task, while also being capable of resisting a variety of popular backdoor attacks and adaptive attacks. MeCa does not require any auxiliary clean data, offering a novel paradigm for backdoor defense and robust training.
@@
  • We evaluate MeCa through extensive experiments on four datasets, which demonstrate a better backdoor defense performance compared to seven SOTA backdoor defense methods. Experimental results validate that MeCa could decrease the attack success rate (average ASR nearly 0.00%) while only incurring negligible accuracy loss. Furthermore, MeCa also demonstrates robust generalization across various model architectures and poisoning ratios.
2 Related Works
@@
In this section, we review the relevant work on backdoor attacks and introduce the development of backdoor defense.
2.1 Backdoor Attack Methods
@@
In recent years, many backdoor attack approaches have been proposed, including patch-based attacks, visibility of trigger attacks, label consistency, and so on. Gu et al.[12] proposed a typical backdoor attack method by adding a specific patch on the samples and modifying the corresponding label as the targeted label. This attack achieves high attack success and has a low-performance impact on the main task. Chen et al.[25] proposed a backdoor attack in which the trigger is blending background rather than a single pixel. This means that the trigger is hard for human beings to notice. Liu et al.[26] proposed a more stealthy backdoor attack that plants reflections as a trigger into the victim model. This attack can be resistant to many existing defense methods. Liu et al.[27] proposed a lightweight backdoor attack that just needs to inverse the neural network to generate a general trojan trigger and fine-tune some layers of the network to implant the trigger. However, all the above backdoor attacks require poisoning the label, which is easier to detect. Therefore, some researchers have also proposed clean-label backdoor attacks. For example, Shafahi et al.[28] proposed a targeted clean-label poisoning attack. This attack crafts poison images that collide with a target image in feature space, thus making it undistinguishable from a network. Because the attacker does not need to control the label, it is more stealthy to conduct a backdoor attack. Turner et al.[29] proposed a clean-label backdoor attack based on adversarial examples and GAN-generated data. The key feature of this attack is that the poisoned samples appear to be consistent with their label and thus seem clean even from human inspection. Chen et al.[8] proposed an Invisible Poisoning Attack (IPA), which is difficult to detect by existing defense methods. This attack not only employs highly stealthy poison training examples with the clean labels (perceptually similar to their clean samples), but also does not need to modify the labels. Li et al.[30] proposed two stealthy backdoor attacks in which the triggers are derived from the covert features. Compared with the existing backdoor attacks, the trigger patterns of this attack method are invisible to human eyes. Moreover, it is difficult to recover the backdoor trigger through the optimization algorithm. Zhu et al.[31] proposed a transferable clean-label poisoning attack in which poison samples are fabricated to surround the targeted sample in feature space. Saha et al.[32] proposed a novel backdoor attack in which the poisoned samples are similar to the clean samples with the correct labels. [33] and [34] proposed defense-resistant backdoor attacks in an outsourced cloud environment.
2.2 Backdoor Defense Methods
@@
Although many backdoor defense methods have been proposed, they have some limitations (e.g., requiring an auxiliary clean dataset) used in practical scenarios. Unlike existing backdoor defense methods, we first explore the relationship between the perturbation and the backdoor. Then, based on the exploration results, we plan to propose a robust backdoor defense method that does not require an auxiliary clean dataset and has a stable performance against various backdoor attacks on different models and poisoning ratios.
3 Preliminaries and Threat Model
In this section, we introduce the mainly related technologies and the threat model in this paper.
3.1 Backdoor Attack
A backdoor attack is to inject a trigger into a model by poisoning the training dataset. During the inference period, the backdoored model performs well on the original task, but outputs specific attacker-chosen responses when the input contains a specific trigger. For more clarity, we formalize the most common backdoor attack method BadNets [12] as follows:
@@ 3.1 Backdoor Attack
where $\times$ is pixel-wise multiplication and $\boldsymbol{x}^{\prime}$ is the poisoned training sample. After modifying the training sample, the corresponding label is fixed with a target label. A successful backdoor model will maintain high accuracy in the main task while outputting the target label when the trigger $\boldsymbol{t}$ appears. After BadNets [12], many different attack variations have been proposed to enhance effectiveness and stealthiness. For example, Blend [25] and PhysicalBA [46] have designed more complex patterns. WaNet [47] has proposed a stealthier input-specific-trigger attack. SIG [48] has proposed a clean-label attack that is stealthier as it does not change the labels. To evade the existing defense method, some adaptive attacks (e.g., TaCT [43], AdaptiveBlend [49] and AdaptivePatch [49]) have also been proposed to improve the attack performance.
Figure 2: Threat Model of MeCa.
3.2 Threat Model
In the scenario considered in MeCa, the model trainer (Defender) intends to train a model on the training data obtained from a third party. A malicious data provider (Attacker) endeavors to manipulate the dataset in order to induce the defender to train a model with a backdoor, thereby facilitating subsequent hijacking of the model. The threat model is depcited in Fig.2. Therefore, in this paper, the defender utilizes MeCa to train a clean model on the poisoned dataset, free from the attacker’s backdoor triggers. Considering the deployment and development of MeCa in real-world scenarios, we provide a detailed description of the goals, knowledge, and capabilities of both the Defender and the Attacker in the following.
@@ 3.2 Threat Model
Figure 3: Framework of MeCa.
4 Design of MeCa
@@
4.1 Overview of MeCa
In this paper, we focus on how to identify the poisoned samples and train a clean model on the poisoned dataset. It is not trivial to satisfy this goal since even a tiny number of poisoned samples can accomplish the backdoor injection process. To address this problem, we first explore the inherent relationship between the perturbation and the backdoor attack theoretically and experimentally. Based on the results of the investigation, we propose a three-stage backdoor defense method dubbed MeCa and the overall framework is shown in Fig.3. There are three main phases in the proposed MeCa. Firstly, we partition a tiny number of clean samples using the difference between clean samples and poisoned samples when subjected to perturbations. An enhanced backdoor model is trained by unlearning the chosen clean samples and relearning the remaining dataset. Then, we accurately identify the poisoned and clean samples with a high degree of precision by leveraging the enhanced backdoor model. Specifically, samples that are misclassified by the enhanced backdoor model are deemed clean, while the remaining samples are classified as poisoned. Subsequently, a clean model can be trained using the clean samples identified. Finally, we relabel the poisoned samples and merge them with the clean ones to obtain a thoroughly clean dataset, enabling the fine-tuning of the clean model for enhanced performance. The design of MeCa is presented in detail as follows.
4.2 Detailed Design of MeCa
@@
4.2.1 Exploring the Relationship between Perturbation and Backdoor
To accurately identify poisoned samples, we need to enhance the backdoor model by unlearning certain clean samples. It is crucial to select a subset of clean samples without relying on an auxiliary dataset and to understand how poisoned samples affect the predictions made by backdoor models. Based on this understanding, we can design metrics to distinguish between clean and poisoned samples.
Our investigation begins with a review of the backdoor attack process, which is similar with [50]. Let $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ represent a clean training set, and let $C:\mathcal{X}\rightarrow\mathcal{Y}$ denote the functionality of the target neural network. Each image $\boldsymbol{x}_{i}$ in $\mathcal{D}$ satisfies $\boldsymbol{x}_{i}\in\mathcal{X}=[0,1]^{C\times W\times H}$, and its corresponding label $y_{i}\in\mathcal{Y}=\{1,\ldots,K\}$, where $K$ is the total number of classes. To initiate an attack, adversaries poison a selection of clean samples $\mathcal{D}_{p}$ using a covert transformation $T(\cdot)$. These poisoned samples are then combined with the clean dataset prior to training the compromised model. This process can be formalized as $\mathcal{D}_{t}=\mathcal{D}\cup\mathcal{D}_{p}$, where $\mathcal{D}_{p}=\{(x_{i}^{\prime},y_{t})|x_{i}^{\prime}=T(x_{i}),(x_{i},y_{i})\in\mathcal{D}_{p}\}$. Various methods have been devised to make $T(\cdot)$ more subtle and harder to detect, or to reduce the assumptions about the adversary’s capabilities, such as the poison ratio $pr=|\mathcal{D}_{p}|/|\mathcal{D}|$. Despite these variations, the ultimate goal remains the same [50]: manipulating the neural network by training a malicious model: